OpenStudy (roadjester):

integrate x*e^x/sqrt(1+e^x)

OpenStudy (anonymous):

Make the substitution that u=sqrt(1+e^x) and you should be able to integrate it :)

OpenStudy (anonymous):

oh wait sorry I didn't see the x in front there

OpenStudy (roadjester):

\[\int\limits_{?}^{?} {xe ^{x}\over \sqrt{1+e^x}}dx\]]

OpenStudy (anonymous):

integrate by parts with u = x and dv/dx = e^x/sqrt(1+e^x) you then do the above substitution to find v

OpenStudy (roadjester):

@eigenschmeigen Final answer: \[2(x-2)\sqrt{1+e ^{x}}+2{\ln {\sqrt{{1+e ^{x}}}+1\over \sqrt{1+e ^{x}}-1}}+C\] When I did the substitution you suggested, I didn't get the answer in the book.

OpenStudy (anonymous):

\[\int\limits{\frac{e^x}{\sqrt{1+e^x}}}dx \] let \[u = 1+e^x\] \[\frac{du}{dx} = e^x\] \[dx= \frac{du}{e^x}\] \[\int\limits\limits{\frac{e^x}{\sqrt{1+e^x}}}dx = \int\limits\limits{\frac{e^x}{\sqrt{u}}}\frac{du}{e^x}\]\[=\int\limits\limits\limits{\frac{1}{\sqrt{u}}}du\]

OpenStudy (anonymous):

i dont really have time to finish, sorry

OpenStudy (roadjester):

@eigenschmeigen That doesn't make any sense! That integral would give me:\[2\sqrt{1+e ^{x}}+C\]

myininaya (myininaya):

This was his plan for you: \[\int\limits_{}^{} x \cdot \frac{e^x}{\sqrt{1+e^x}} dx= x \cdot 2 \sqrt{1+e^x}-\int\limits_{}^{} (x)' 2 \sqrt{1+e^x} dx\]

myininaya (myininaya):

hint for that one part: http://openstudy.com/study#/updates/4f7b211ae4b0884e12aafc4b

myininaya (myininaya):

Ok I will be back on tomorrow if more help is needed peace

OpenStudy (roadjester):

@myininaya I'm not sure where your method progresses to, but an integration by parts plus a trig substitution is too complex. No offense.

OpenStudy (apoorvk):

I think roadjester's method is a bit easier, but then, myininaya's way is another new thing I learnt today. :)

OpenStudy (apoorvk):

Sorry, I meant eigenschmeigen's instead of roadjester's.

OpenStudy (anonymous):

yeah, by parts like @myininaya has done, then it requires a trig substitution